# Complexity of Local Search for Euclidean Clustering Problems

### v1

@@ 2.3 Strategy
Both Squared Euclidean Max Cut and $k$-Means are NP-hard [ageevComplexityWeightedMaxcut2014, ADHP09]. The reductions used to prove this are quite similar, and can be straightforwardly adapted into PLS-reductions: In the case of Squared Euclidean Max Cut/Flip, we obtain a reduction from Odd Min Bisection/Flip, while for $k$-Means/Hartigan–Wong we obtain a reduction from Densest Cut/Flip. In fact, the latter reduction even works for $k=2$. These results are given in LABEL:lemma:densestcut_to_hartigan ($k$-Means), and LABEL:lemma:bisection_to_emaxcut_nosquare and LABEL:lemma:bisection_to_emaxcut ((Squared) Euclidean Max Cut).
What remains then is to show that the problems we reduce from are also PLS-complete, which will take up the bulk of this paper. Figure 1 shows the reduction paths we use.
The starting point will be the PLS-completeness of Max Cut-$5$/Flip, which was shown by Elsässer and Tscheuschner [elsasserSettlingComplexityLocal2011]. An obvious next step might then be to reduce from this problem to Max Bisection/Swap, and then further reduce to Odd Max Bisection/Flip. Unfortunately, this turns out to be rather difficult, as the extra power afforded by the Swap neighborhood is not so easily reduced back to the Flip neighborhood. This makes it rather challenging to obtain a map from feasible solutions of Odd Max Bisection/Flip to Max Bisection/Swap that preserves local optimality.
We thus take a detour through Odd Half Pos NAE $3$-SAT/Flip in Lemma 3.3, which then reduces down to Odd Half Pos NAE $2$-SAT/Flip in Lemma 3.5 and finally to Odd Max Bisection/Flip in Lemma 3.9, using a reduction by Schäffer and Yannakakis [schafferSimpleLocalSearch1991].
From this point, hardness of Squared Euclidean Max Cut/Flip (and with a little extra work, Euclidean Max Cut/Flip) is easily obtained. For $k$-Means/Hartigan–Wong, we need some more effort, as we still need to show hardness of Densest Cut/Flip. Luckily, this can be done by reducing from Odd Max Bisection/Flip as well, as proved in Lemma 4.1.
@@ 3 Reduction to Odd Min/Max Bisection
The goal of this section is to obtain PLS-completeness of Odd Min/Max Bisection/Flip, from which we can reduce further to our target problems; see Figure 1. We will first construct a PLS-reduction from Max Cut-$5$/Flip to Odd Half Pos NAE $3$-SAT/Flip in Lemma 3.3.
@@
Lemma 3.1.
Distinct Max Cut-$5$/Flip is PLS-complete.
Proof 3.2.
@@
Correctness. Let $s$ and $s^{\prime}$ be two neighboring solutions to $I^{\prime}$. The neighbor $s^{\prime}$ is obtained by flipping some vertex. Observe that the only edge weights that enter $c(s^{\prime},I)-c(s,I)$ are those incident to this vertex. Thus, since every vertex of $H$ has odd degree, we have an odd number of terms. Since every term is increased by exactly $1$ in $I^{\prime}$, if $c(s^{\prime},I)=c(s,I)$, then $c(s^{\prime},I^{\prime})\neq c(s,I^{\prime})$. Moreover, if $c(s^{\prime},I)\neq c(s,I)$, then the two differ by at least 10, while the value of $c(s^{\prime},I^{\prime})-c(s,I^{\prime})$ differs from $c(s^{\prime},I)-c(s,I)$ by at most 5. Hence, $c(s,I)\neq c(s^{\prime},I)$ implies $c(s,I^{\prime})\neq c(s^{\prime},I^{\prime})$, and so any two neighboring solutions to $I^{\prime}$ have distinct cost.
We claim that a locally optimal solution to $I^{\prime}$ is also locally optimal for $I$. We prove the contrapositive. Suppose a solution $s$ to $I$ is not locally optimal, so there exists a neighbor $s^{\prime}$ of $s$ such that $c(s^{\prime},I)-c(s,I)\geq 10$. Following the same argument as in the previous paragraph, $c(s^{\prime},I^{\prime})-c(s,I^{\prime})$ differs from $c(s^{\prime},I)-c(s,I)$ by at most $5$. Hence, $c(s^{\prime},I^{\prime})-c(s,I^{\prime})\geq 5$. Therefore, $s$ is also not locally optimal for $I^{\prime}$.
Note that the reduction used in Lemma 3.1 is not tight. Hence, to prove the last two items of Theorem 1.3, simply applying the reductions from Figure 1 is not sufficient, as these properties (viz. PSPACE-completeness of the standard algorithm problem and the existence of certain hard instances) do not necessarily hold for Distinct Max Cut-$5$/Flip. We must instead prove that they hold for this problem separately. To accomplish this, we recall a construction by Monien and Tscheuschner [monienPowerNodesDegree2010] that shows these properties for Max Cut-$4$/Flip. It can be verified straightforwardly that the construction they use is already an instance of Distinct Max Cut-$4$/Flip. We omit the proof for space reasons.
In the remainder of this work, we obtain a sequence of tight reductions starting from Distinct Max Cut-$5$/Flip to all of our other considered problems. First, we reduce from Distinct Max Cut-$5$/Flip to Odd Half Pos NAE $3$-SAT/Flip.
[…]
Figure 2: Schematic overview of the reduction used in the proof of Lemma 3.3. On the left we have a vertex $v\in V$ and its neighbors $\{u_{1},u_{2},u_{3}\}$ in a Max Cut instance, with weights on the edges between $v$ and its neighbors. The $\operatorname{NAE}$ clauses on the right are the clauses constructed from $v$. In the actual reduction, these clauses are added for all level 3 variables. The right-most column shows the weights assigned to the clauses. The middle column shows for the level 3 clauses which subset of $N(v)$ corresponds to which clause. The constants $L$ and $M$ are chosen so that $1\ll L\ll M$.
Lemma 3.3.
Next, we briefly sketch and motivate some of the ideas in the reduction. See also Figure 2.
[…]
Sketch of Proof for Lemma 3.3. The first idea is to embed the Distinct Max Cut-5 instance, given by a weighted graph $G=(V,E)$, in Pos NAE SAT. This can be done rather straightforwardly, by a reduction used by Schäffer and Yannakakis [schafferSimpleLocalSearch1991]: each vertex becomes a variable, and an edge $uv$ becomes a clause $\operatorname{NAE}(u,v)$. This instance is directly equivalent to the original Distinct Max Cut-5 instance. We call these variables the level 1 variables, and the clauses the level 1 clauses. A level 1 clause $\operatorname{NAE}(u,v)$ gets a weight $M\cdot w(uv)$ for some large integer $M$.
@@
Given the reduction so far, suppose we have some locally optimal feasible truth assignment $s$. We partition the variables into the sets $T$ and $F$ of true and false variables; thus, $(T,F)$ is the cut induced by $s$. Suppose $|T|=|F|+1$. If no level 1 variable can be flipped from $T$ to $F$, then also no vertex can flip from $T$ to $F$ in the induced cut. However, we run into trouble when there exists some $v\in F$ that can flip in the cut. Since $|F|<|T|$, we are not allowed to flip the level 1 variable $v$, and so the truth assignment may be locally optimal even though the induced cut is not.
@@
[…]
Correctness. It now remains to show that a locally optimal truth assignment induces a locally optimal cut. In the following, given a truth assignment $s$, we will denote by $T(s)$ the set of variables assigned true by $s$ and with $F(s)$ the set assigned false by $s$. If the truth assignment $s$ is clear from the context, we may simply write $T$ and $F$.
@@
Thus, $v$ can flip from true to false in $s$, proving the claim.
Let $s$ be a locally optimal truth assignment for $S$. For the remainder of the proof, we assume that $|T(s)|=|F(s)|+1$; this is w.l.o.g., since the problem is symmetrical under inversion of the truth value of all variables simultaneously. Since $s$ is locally optimal, no level 1 variable in $T$ can be flipped for a positive gain. Hence, no vertex in the corresponding Max Cut-5 solution can be flipped from $T\cap V$ to $F\cap V$ by Claim 1.
@@
We know from Claim 1 that $v$ is unable to flip in the cut induced from $s$, as otherwise $s$ would not be locally optimal. Thus,
@@
We see that flipping $q_{v}$ would strictly increase the cost, contradicting local optimality of $s$. Thus, we must have $q_{v}\in F$, which then implies that flipping $a_{i}$ from true to false does not change the status of $C_{i}(v,Q)$. The claim follows.
@@
We will show that we must have $q_{v}\in T$. Suppose towards a contradiction that $q_{v}\in F$. Then flipping $a_{i}\in T$ unsatisfies $C_{i}(v,Q)$, yielding a positive contribution to the cost. Moreover, by Claim 2, flipping $a_{i}$ cannot yield any negative contributions to the cost, and hence this flip is improving. This contradicts local optimality of $s$, and thus we must have $q_{v}\in T$.
@@
as in the proof of Claim 2. But the quantity in brackets is exactly the change in cost that would be incurred by flipping $v$ in the induced cut, which was assumed improving; hence, flipping $q_{v}$ would increase the cost by at least $L-2^{5}\cdot N=2^{6}N-2^{5}N>0$, contradicting local optimality of $s$. Thus, no $v\in V\cap F$ can be eligible to flip in the induced cut.
Together with Claim 1, we can conclude that if $s$ is locally optimal, then there exists no vertex $v\in V$ that can flip in the cut induced by $s$. Thus, local optima for Odd Half Pos NAE $3$-SAT/Flip map to a local optima of Max Cut-$5$/Flip, satisfying the essential property of PLS-reductions. Moreover, the reduction can be achieved in polynomial time, and so it is a valid PLS-reduction.
[…]
Tightness. Let $x$ be an instance of Distinct Max Cut-5 and let $(f,g)$ be the reduction given above. Let $\mathcal{R}$ be the entire set of feasible truth assignments for $f(x)$. This obviously contains all local optima of $f(x)$, satisfying property 1 of tight reductions. Given some cut $(A,B)$ in $G$, we can construct $s\in\mathcal{R}$ by setting the variables in $A\cap S$ to true and those in $B\cap S$ to false, assigning the level 2 variables a truth value arbitrarily, and assigning the level 3 variables such that $|T|=|F|+1$. This satisfies property 2 of tight reductions. For property 3, suppose the transition graph of the reduced instance $f(x)$ contains a directed path $P$ from $s\in\mathcal{R}$ to $s^{\prime}\in\mathcal{R}$ with all internal vertices outside of $\mathcal{R}$. Since $\mathcal{R}$ is the entire set of feasible truth assignments, this implies that $P$ consists of the single arc $ss^{\prime}$. If $s$ and $s^{\prime}$ differ in the flip of a level 2 or level 3 variable, then $g(s,x)=g(s^{\prime},x)$, and we are done. Suppose then that $s$ and $s^{\prime}$ differ in the flip of a level 1 variable $v$. To conclude, we must show that the transition graph of $x$ contains an arc from $g(s,x)$ to $g(s^{\prime},x)$. Suppose to the contrary that no such arc exists. This implies that flipping the vertex $v$ in $g(s,x)$ reduces the cut value by at least $1$. Thus, flipping the variable $v$ in $s$ reduces the cost of the truth assignment by at least $M-5\Delta_{\max{}}L-2^{5}N>0$, which contradicts the existence of the arc $ss^{\prime}$.
A reduction from Pos NAE $3$-SAT/Flip to Max Cut/Flip was provided by Schäffer and Yannakakis [schafferSimpleLocalSearch1991]. Since Max Cut is equivalent to Pos NAE 2-Sat, we can use the same reduction to reduce from Odd Half Pos NAE $3$-SAT/Flip to Odd Half Pos NAE $2$-SAT/Flip. We write the proof here for the sake of completeness.
Lemma 3.5 (Schäffer and Yannakakis [schafferSimpleLocalSearch1991]).
@@
If the weight of the original clause is $W$, then the resulting clauses each get a weight of $W/2$. Note that this may result in several clauses consisting of the same pair of variables. For the purposes of this analysis, we consider each of these clauses to be distinct; however, one could equivalently merge these clauses into one clause with weight equal to the the total weight of the merged clauses.
Let $s$ be some feasible solution. If a clause $C$ of size three of weight $W$ in $I$ is unsatisfied under $s$, then its resulting clauses in $I^{\prime}$ are also unsatisfied. If $C$ is satisfied under $s$, then exactly two of its three resulting clauses are satisfied, yielding a contribution of $W$ to the objective function. Thus, the cost of a solution $s$ in $I$ is exactly the cost of $s$ in $I^{\prime}$. The reduction clearly preserves local optima, and is tight since the transition graphs of the two instances are identical.
While our reductions so far have used negative-weight clauses in Odd Half Pos NAE $k$-SAT/Flip, it may be of interest to have a PLS-completeness result also when all clauses have non-negative weight. We briefly show that this can be done, as a corollary of Lemma 3.5.
Corollary 3.7.
Odd Half Pos NAE $2$-SAT/Flip is PLS-complete even when all clauses have non-negative weight. Moreover, the reduction to this restricted version of the problem is tight.
Proof 3.8.
@@
which is simply the weight of the same assignment in the original instance, plus a fixed term. Thus, the local optima of the reduced instance are identical to those of the original instance. Tightness follows since the transition graphs of the two instances are identical.
@@
Proof 3.10.
A reduction from Pos NAE $2$-SAT/Flip to Max Cut/Flip is given by Schäffer and Yannakakis [schafferSimpleLocalSearch1991]; it is essentially the reverse direction of the reduction given at the start of the proof of Lemma 3.3. It is easy to see that this reduction also works with our constraint on the number of true and false variables, which yields a reduction to Odd Max Bisection/Flip. By Corollary 3.7, we can assume that the clauses in the original instance all have non-negative weight, and thus this also holds for the edges in the resulting instance of Odd Max Bisection.
@@
We take the resulting instance as an instance of Odd Min Bisection/Flip. The local maxima of the Odd Max Bisection/Flip instance become local minima, and so locally optimal solutions are preserved, concluding the proof. For tightness, we observe that the transition graphs of the instances are identical.

### v3

@@ 2.3 Strategy
Both Squared Euclidean Max Cut and $k$-Means are NP-hard [2, 3]. The reductions used to prove this are quite similar, and can be straightforwardly adapted into PLS-reductions: In the case of Squared Euclidean Max Cut/Flip, we obtain a reduction from Odd Min Bisection/Flip, while for $k$-Means/Flip we obtain a reduction from Densest Cut/Flip. The latter reduction even works for $k=2$. These results are given in Lemma 12 ($k$-Means), and Lemmas 15 and 14 ((Squared) Euclidean Max Cut).
What remains then is to show that the problems we reduce from are also PLS-complete, which takes up the bulk of the work. Figure 1 shows the reduction paths we use.
The starting point will be the PLS-completeness of Max Cut-$5$/Flip, which was shown by Elsässer and Tscheuschner [16]. An obvious next step might then be to reduce from this problem to Max Bisection/Swap, and then further reduce to Odd Max Bisection/Flip. Unfortunately, this turns out to be rather difficult, as the extra power afforded by the Swap neighborhood is not so easily reduced back to the Flip neighborhood. Using this strategy, we can only obtain PLS-hardness of the 2-Flip neighborhood for Squared Euclidean Max Cut, where two points may flip in a single iteration.
We thus take a detour through Odd Half Pos NAE $3$-SAT/Flip in Lemma 6, which then reduces down to Odd Half Pos NAE $2$-SAT/Flip in Lemma 7 and finally to Odd Max Bisection/Flip in Lemma 9, using a reduction by Schäffer and Yannakakis [42].
From this point, hardness of Squared Euclidean Max Cut/Flip (and with a little extra work, Euclidean Max Cut/Flip) is easily obtained. For $k$-Means/Flip, we need some more effort, as we still need to show hardness of Densest Cut/Flip. Luckily, this can be done by reducing from Odd Max Bisection/Flip as well, as proved in Lemma 10.
@@ 3 Reduction to Odd Min/Max Bisection
The goal of this section is to obtain PLS-completeness of Odd Min/Max Bisection/Flip, from which we can reduce further to our target problems; see Figure 1. We will first construct a PLS-reduction from Max Cut-$5$/Flip to Odd Half Pos NAE $3$-SAT/Flip in Lemma 6.
@@
Lemma 4.
Distinct Max Cut-$5$/Flip is PLS-complete. More precisely, there exists a PLS-reduction from Max Cut-$5$/Flip to Distinct Max Cut-$5$/Flip.
Proof.
@@
Correctness.
Let $s$ and $s^{\prime}$ be two neighboring solutions to $I^{\prime}$. The neighbor $s^{\prime}$ is obtained by flipping some vertex. Observe that the only edge weights that enter $c(s^{\prime},I)-c(s,I)$ are those incident to this vertex. Thus, since every vertex of $H$ has odd degree, we have an odd number of terms. Since every term is changed by exactly $1$ in $I^{\prime}$, if $c(s^{\prime},I)=c(s,I)$, then $c(s^{\prime},I^{\prime})\neq c(s,I^{\prime})$. Moreover, if $c(s^{\prime},I)\neq c(s,I)$, then the two differ by at least 10, while the value of $c(s^{\prime},I^{\prime})-c(s,I^{\prime})$ differs from $c(s^{\prime},I)-c(s,I)$ by at most 5. Hence, $c(s,I)\neq c(s^{\prime},I)$ implies $c(s,I^{\prime})\neq c(s^{\prime},I^{\prime})$, and so any two neighboring solutions to $I^{\prime}$ have distinct cost.
We claim that a locally optimal solution to $I^{\prime}$ is also locally optimal for $I$. We prove the contrapositive. Suppose a solution $s$ to $I$ is not locally optimal, so there exists a neighbor $s^{\prime}$ of $s$ such that $c(s^{\prime},I)-c(s,I)\geq 10$. Following the same argument as in the previous paragraph, $c(s^{\prime},I^{\prime})-c(s,I^{\prime})$ differs from $c(s^{\prime},I)-c(s,I)$ by at most $5$. Hence, $c(s^{\prime},I^{\prime})-c(s,I^{\prime})\geq 5$. Therefore, $s$ is also not locally optimal for $I^{\prime}$. ∎
Note that the reduction used in Lemma 4 is not tight. Hence, to prove the last two items of Theorem 3, simply applying the reductions from Figure 1 is not sufficient, as these properties (viz. PSPACE-completeness of the standard algorithm problem and the existence of certain hard instances) do not necessarily hold for Distinct Max Cut-$5$/Flip. We must instead separately prove that they hold for this problem. To accomplish this, we recall a construction by Monien and Tscheuschner [38] that shows these properties for Max Cut-$4$/Flip. It can be verified straightforwardly that the construction they use is already an instance of Distinct Max Cut-$4$/Flip, as we show.
To verify this claim, we first state some terminology used by Monien and Tscheuschner. Let $G=(V,E)$ be a graph of maximum degree four. Let $u\in V$, and let $\{a_{u},b_{u},c_{u},d_{u}\}$ be the edges incident to $u$; if $u$ has degree less than four, we consider the weights of the non-existent edges to be zero. Assume in the following that $w(a_{u})\geq w(b_{u})\geq w(c_{u})\geq w(c_{u})$. We distinguish three types of vertices. A vertex $u$ is of:
  1. Type I,
    if $w(a_{u})>w(b_{u})+w(c_{u})+w(d_{u})$;
  2. Type II,
    if $w(a_{u})+w(d_{u})>w(b_{u})+w(c_{u})$ and $w(a_{u})<w(b_{u})+w(c_{u})+w(d_{u})$;
  3. Type III,
    if $w(a_{u})+w(d_{u})<w(b_{u})+w(c_{u})$.
While general weighted graphs may contain vertices not of these three types, the graphs constructed by Monien and Tscheuschner only contain Type I, II and III vertices.
Given a cut of $V$, a vertex $u$ is called happy if flipping $u$ cannot increase the weight of the cut. Monien and Tscheuschner note that a vertex of type I is happy if and only if the edge $a_{u}$ is in the cut, a vertex of type II is happy if and only if $a_{u}$ and at least one of $\{b_{u},c_{u},d_{u}\}$ is in the cut, and a vertex of type III is happy if and only if at least two of $\{a_{u},b_{u},c_{u}\}$ are in the cut.
Lemma 5.
Let the graph above define an instance $x$ of Max Cut-$4$/Flip, and let $T(x)$ be its transition graph. Suppose $s$ and $s^{\prime}$ are neighbors in $T(x)$ which differ by the flip of a vertex of type I, type II, or type III. Then $c(s,x)\neq c(s^{\prime},x)$.
Proof.
Let $u$ be the vertex in which $s$ and $s^{\prime}$ differ, with incident edges $\{a_{u},b_{u},c_{u},d_{u}\}$. Assume the weights of these edges are ordered as $w(a_{u})\geq w(b_{u})\geq w(c_{u})\geq w(d_{u})$. Furthermore, we assume w.l.o.g. that $u$ is happy in $s$. We consider the three possibilities for the type of $u$.
Type I.
If $u$ is happy in $s$, then $a_{u}$ is in the cut. Thus, we have $c(s,x)-c(s^{\prime},x)\geq w(a_{u})-(w(b_{u})+w(c_{u})+w(d_{u}))>0$. The claim follows.
Type II.
Since $u$ is happy in $s$, we know that $a_{u}$ and at least one of $\{b_{u},c_{u},d_{u}\}$ are in the cut in $s$. Assume that $d_{u}$ is in the cut with $a_{u}$. Then $c(s,x)-c(s^{\prime},x)\geq w(a_{u})+w(d_{u})-(w(b_{u})+w(c_{u}))>0$. Now observe that if $d_{u}$ is swapped out for $b_{u}$ or $c_{u}$, then $c(s,x)-c(s^{\prime},x)$ cannot decrease, since $w(d_{u})\leq w(c_{u})\leq w(b_{u})$ by assumption. Thus, flipping $u$ changes the weight of the cut if $u$ is of type II.
Type III.
Since $u$ is happy in $s$, at least two of $\{a_{u},b_{u},c_{u}\}$ are in the cut. We distinguish two cases.
Case 1.
Assume first that $a_{u}$ is not in the cut, so that $b_{u}$ and $c_{u}$ are both in the cut. Then $c(s,x)-c(s^{\prime},x)\geq w(b_{u})+w(c_{u})-(w(a_{u})+w(d_{u}))>0$.
Case 2.
Now we assume that $a_{u}$ is in the cut. Then at least one of $b_{u}$ and $c_{u}$ is in the cut with $a_{u}$. Thus, since $w(a_{u})\geq w(b_{u})\geq w(c_{u})$, we have $c(s,x)-c(s^{\prime},x)\geq w(a_{u})+w(c_{u})-(w(b_{u})+w(d_{u}))\geq w(b_{u})+w(c_{u})-(w(a_{u})+w(d_{u}))>0$. Again, flipping $u$ strictly changes the weight of the cut.
This covers all possibilities, and so we conclude that flipping a vertex of any of the three types defined above strictly changes the weight of the cut. ∎
Since the graphs constructed by Monien and Tscheuschner only contain vertices of these three types, we can conclude that these graphs are instances of Distinct Max Cut-4 and thus also of Distinct Max Cut-5.
In the remainder of Section 3 and subsequently in Section 4 we will obtain the PLS-reductions shown in Figure 1. Since the reductions from Distinct Max Cut-$5$/Flip are tight, and by the properties of tight PLS-reductions and the results of Monien and Tscheuschner, Theorem 3 then follows.
In the remainder of this work, we present a sequence of tight reductions starting from Distinct Max Cut-$5$/Flip to all of our other considered problems. First, we reduce from Distinct Max Cut-$5$/Flip to Odd Half Pos NAE $3$-SAT/Flip.
Lemma 6.
As mentioned in Section 2.3, it may seem more straightforward to reduce from Min Bisection/Swap to Odd Min Bisection/Flip. The problem with this approach is that a solution to Odd Min Bisection/Flip can be locally optimal for two reasons: either no vertex can flip to obtain a cut of larger weight, or the vertices that could are in the smaller part of the partition. This makes the Flip neighborhood much less powerful than Swap in this problem variant; we were thus not able to find a direct reduction from Min Bisection/Swap. Instead, we apply a new technique that allows us to prove PLS-hardness for this very restricted problem.
We first prove PLS-hardness of Odd Half Pos NAE $3$-SAT/Flip, and subsequently use existing reductions to obtain hardness of Odd Min Bisection/Flip. With the expressiveness of this SAT variant we gain a great deal of freedom to handle the problem restrictions. The main challenge is in encoding the restriction that the number of true and false variables must differ by exactly one without weakening the neighborhood.
We achieve this in two steps. First, by adding a large number of extra variables, we can embed any solution to our original instance of Max Cut-5 into Odd Half Pos NAE 3-SAT while still satisfying the size constraints. Second, we must restore the power of the Flip neighborhood in this instance. By adding new clauses and variables to the instance, we simulate a flip in the original Max Cut-5 instance by a sequence of at most two flips in its image instance. This requires carefully constructing clauses which encode subsets of neighborhoods of vertices in the original Max Cut-5 instance. Since each vertex in this instance has bounded degree, there are a constant number of subsets of the neighborhood of any vertex, yielding a polynomial-time reduction.
As far as we are aware, this technique for overcoming size constraints in local search problems is novel. We believe that it may be useful to prove PLS-hardness results for simple heuristics for other size-constrained problems, such as balanced clustering problems.
Next, we briefly sketch and motivate some of the ideas in the reduction in more detail, before proceeding to the proof. See also Figure 2.
@@ 3. Type III,
Figure 2: Schematic overview of the reduction used in the proof of Lemma 6. On the left we have a vertex $v\in V$ and its neighbors $\{u_{1},u_{2},u_{3}\}$ in a Max Cut instance, with weights on the edges between $v$ and its neighbors. The $\operatorname{NAE}$ clauses on the right are the clauses constructed from $v$. In the actual reduction, these clauses are added for all level 3 variables. The right-most column shows the weights assigned to the clauses. The middle column shows for the level 3 clauses which subset of $N(v)$ corresponds to which clause. The constants $L$ and $M$ are chosen so that $1\ll L\ll M$.
Sketch of Proof for Lemma 6.
We first embed the Distinct Max Cut-5 instance, given by a weighted graph $G=(V,E)$, in Pos NAE SAT. This can be done rather straightforwardly, by a reduction used by Schäffer and Yannakakis [42]: each vertex becomes a variable, and an edge $uv$ becomes a clause $\operatorname{NAE}(u,v)$. This instance is directly equivalent to the original Distinct Max Cut-5 instance. We call these variables the level 1 variables, and the clauses the level 1 clauses. A level 1 clause $\operatorname{NAE}(u,v)$ gets a weight $M\cdot w(uv)$ for some large integer $M$.
@@
Given the reduction so far, suppose we have some locally optimal feasible truth assignment $s$. We partition the variables into the sets $T$ and $F$ of true and false variables; thus, $(T,F)$ is the cut induced by $s$. Suppose $|T|=|F|+1$. If no level 1 variable can flip from $T$ to $F$, then also no vertex can flip from $T$ to $F$ in the induced cut. However, we run into trouble when there exists some $v\in F$ that can flip in the cut. Since $|F|<|T|$, we are not allowed to flip the level 1 variable $v$, and so the truth assignment may be locally optimal even though the induced cut is not.
@@
Correctness.
It now remains to show that a locally optimal truth assignment induces a locally optimal cut. In the following, given a truth assignment $s$, we will denote by $T(s)$ the set of variables assigned true by $s$ and with $F(s)$ the set assigned false by $s$. If the truth assignment $s$ is clear from the context, we may simply write $T$ and $F$.
@@
Thus, $v$ can flip from true to false in $s$, proving the claim. ∎
Let $s$ be a locally optimal truth assignment for $S$. For the remainder of the proof, we assume that $|T(s)|=|F(s)|+1$; this is w.l.o.g., since the problem is symmetrical under inversion of the truth value of all variables simultaneously. Since $s$ is locally optimal, no level 1 variable in $T$ can be flipped for a positive gain. Hence, no vertex in the corresponding Max Cut-5 solution can be flipped from $T\cap V$ to $F\cap V$ by Claim Claim.
@@
We know from Claim Claim that $v$ is unable to flip in the cut induced from $s$, as otherwise $s$ would not be locally optimal. Thus,
@@
We see that flipping $q_{v}$ would strictly increase the cost, contradicting local optimality of $s$. Thus, we must have $q_{v}\in F$, which then implies that flipping $a_{i}$ from true to false does not change the status of $C_{i}(v,Q)$. The claim follows. ∎
@@
We will show that we must have $q_{v}\in T$. Suppose towards a contradiction that $q_{v}\in F$. Then flipping $a_{i}\in T$ unsatisfies $C_{i}(v,Q)$, yielding a positive contribution to the cost. Moreover, by Claim Claim, flipping $a_{i}$ cannot yield any negative contributions to the cost, and hence this flip is improving. This contradicts local optimality of $s$, and thus we must have $q_{v}\in T$.
@@
as in the proof of Claim Claim. But the quantity in brackets is exactly the change in cost that would be incurred by flipping $v$ in the induced cut, which was assumed improving; hence, flipping $q_{v}$ would increase the cost by at least $L-2^{5}\cdot N=2^{6}N-2^{5}N>0$, contradicting local optimality of $s$. Thus, no $v\in V\cap F$ can be eligible to flip in the induced cut.
Together with Claim Claim, we can conclude that if $s$ is locally optimal, then there exists no vertex $v\in V$ that can flip in the cut induced by $s$. Thus, local optima for Odd Half Pos NAE $3$-SAT/Flip map to a local optima of Max Cut-$5$/Flip, satisfying the essential property of PLS-reductions. Moreover, the reduction can be achieved in polynomial time, and so it is a valid PLS-reduction.
Tightness.
Let $x$ be an instance of Distinct Max Cut-5 and let $(f,g)$ be the reduction given above. Let $\mathcal{R}$ be the entire set of feasible truth assignments for $f(x)$. This obviously contains all local optima of $f(x)$, satisfying property 1 of tight reductions. Given some cut $(A,B)$ in $G$, we can construct $s\in\mathcal{R}$ by setting the variables in $A\cap S$ to true and those in $B\cap S$ to false, assigning the level 2 variables a truth value arbitrarily, and assigning the level 3 variables such that $|T|=|F|+1$. This satisfies property 2 of tight reductions. For property 3, suppose the transition graph of the reduced instance $f(x)$ contains a directed path $P$ from $s\in\mathcal{R}$ to $s^{\prime}\in\mathcal{R}$ with all internal vertices outside of $\mathcal{R}$. Since $\mathcal{R}$ is the entire set of feasible truth assignments, this implies that $P$ consists of the single arc $ss^{\prime}$. If $s$ and $s^{\prime}$ differ in the flip of a level 2 or level 3 variable, then $g(s,x)=g(s^{\prime},x)$, and we are done. Suppose then that $s$ and $s^{\prime}$ differ in the flip of a level 1 variable $v$. To conclude, we must show that the transition graph of $x$ contains an arc from $g(s,x)$ to $g(s^{\prime},x)$. Suppose to the contrary that no such arc exists. This implies that flipping the vertex $v$ in $g(s,x)$ reduces the cut value by at least $1$. Thus, flipping the variable $v$ in $s$ reduces the cost of the truth assignment by at least $M-5\Delta_{\max{}}L-2^{5}N>0$, which contradicts the existence of the arc $ss^{\prime}$. ∎
A reduction from Pos NAE $3$-SAT/Flip to Max Cut/Flip was provided by Schäffer and Yannakakis [42]. Since Max Cut is equivalent to Pos NAE 2-Sat, we can use the same reduction to reduce from Odd Half Pos NAE $3$-SAT/Flip to Odd Half Pos NAE $2$-SAT/Flip. We write the proof here for the sake of completeness.
Lemma 7 (Schäffer and Yannakakis [42]).
@@
If the weight of the original clause is $W$, then the resulting clauses each get a weight of $W/2$. Note that this may result in several clauses consisting of the same pair of variables. For the purposes of this analysis, we consider each of these clauses to be distinct; however, one could equivalently merge these clauses into one clause with weight equal to the total weight of the merged clauses.
Let $s$ be some feasible solution. If a clause $C$ of size three of weight $W$ in $I$ is unsatisfied under $s$, then its resulting clauses in $I^{\prime}$ are also unsatisfied. If $C$ is satisfied under $s$, then exactly two of its three resulting clauses are satisfied, yielding a contribution of $W$ to the objective function. Thus, the cost of a solution $s$ in $I$ is exactly the cost of $s$ in $I^{\prime}$. The reduction clearly preserves local optima, and is tight since the transition graphs of the two instances are identical. ∎
While our reductions so far have used negative-weight clauses in Odd Half Pos NAE $k$-SAT/Flip, it may be of interest to have a PLS-completeness result also when all clauses have non-negative weight. We briefly show that this can be done, as a corollary of Lemma 7.
Corollary 8.
Odd Half Pos NAE $2$-SAT/Flip is PLS-complete even when all clauses have non-negative weight. More precisely, there exists a tight PLS-reduction from Odd Half Pos NAE $2$-SAT/Flip to Odd Half Pos NAE $2$-SAT/Flip where all clauses have non-negative weight.
Proof.
@@
which is simply the weight of the same assignment in the original instance, plus a fixed term. Thus, the local optima of the reduced instance are identical to those of the original instance. Tightness follows since the transition graphs of the two instances are identical. ∎
@@
A reduction from Pos NAE $2$-SAT/Flip to Max Cut/Flip is given by Schäffer and Yannakakis [42]; it is essentially the reverse direction of the reduction given at the start of the proof of Lemma 6. It is easy to see that this reduction also works with our constraint on the number of true and false variables, which yields a reduction to Odd Max Bisection/Flip.
Proof.
By Corollary 8, we can assume that the clauses in the original instance all have non-negative weight, and thus this also holds for the edges in the resulting instance of Odd Max Bisection.
@@
We take the resulting instance as an instance of Odd Min Bisection/Flip. The local maxima of the Odd Max Bisection/Flip instance become local minima, and so locally optimal solutions are preserved, concluding the proof. For tightness, we observe that the transition graphs of the instances are identical. ∎